# MTFM: A Scalable and Alignment-free Foundation Model for Industrial Recommendation in Meituan

### v1

@@ Abstract
1 Introduction
The landscape of Large Language Models (LLM) has been redefined by the era of Foundation Models (FMs), where unified architectures have evolved from unimodal to multimodal (Bai et al., 2025; Yang et al., 2025; Team et al., 2025; Fu et al., 2025; Liu et al., 2026). In parallel, the recommendation community has begun to validate the Scaling Laws (Kaplan et al., 2020; Hoffmann et al., 2022), primarily focusing on the expansion of model parameters and data volume within isolated scenarios (Zhai et al., 2024; Zhu et al., 2025; Zhang et al., 2024a; Wang et al., 2025; Zhou et al., 2025a; Han et al., 2024). However, the true potential of a Recommendation Foundation Model remains untapped as long as models are confined to single-scenario silos. Inspired by the success of multimodal FMs, we argue that the next frontier lies in bridging cross-scenario heterogeneity. By extending from single-scene optimization to a unified multi-scenario framework, we can leverage diverse behavioral signals to further unlock the benefits of Scaling Laws, thereby elevating the performance upper bound and establishing a more scalable paradigm for recommendation.
@@ 1 Introduction
Existing multi-scenario modeling efforts (Sheng et al., 2021a; Shu et al., 2024; Zhang et al., 2024c; Xiaoyu et al., 2025) primarily follow a "harmonize-then-decompose" paradigm. These approaches typically aggregate data from multiple scenarios and enforce input homogeneity with a fixed template. Any heterogeneous components must be either discarded or handled through an additional padding operator. Structurally, they rely on disentangling model parameters into domain-invariant and domain-specific components (e.g., via MoE-based architectures (Ma et al., 2018b; Fedus et al., 2022)) to capture shared knowledge while preserving scenario-specific nuances.
@@
Empirical Validation: Extensive offline and online evaluations at Meituan demonstrate significant gains, proving MTFM's real-world effectiveness.
2 Methodology
@@
2.1 User Sample Aggregation with Multi-Scenario Data
@@
2.2 MTFM Model Architecture
@@ 2.2.1 Heterogeneous Tokenization
2.2.2 Hybrid Target Attention Architecture
@@
The embeddings of the T-tokens from the final layer are input into an MMoE(Ma et al., 2018b) module to compute prediction scores for various objectives across multiple scenarios.
@@
2.3 Training and Deployment Optimization
2.3.1 Data Pipeline of MTFM
@@ 3 Experiments
RQ1: How does the MTFM perform across multiple heterogeneous multi-scenario tasks compared to other baseline models?
RQ2: How does the hybrid attention architecture in MTFM affect training and inference performance?
RQ3: How does the MTFM perform in terms of scalability, including model computational complexity scalability and data scalability?
@@
3.1 Experiments Setup
3.1.1 Datasets
@@
| HP | 18.53B | 1.08B | 176.77M | ||
| PHF | 15.29B | 359.14M | 104.73M | ||
| SQS | 2.24B | 85.34M | 9.92M |
To accommodate the heterogeneous multi-scenario business data proposed in this work, we construct our offline experimental dataset from a real-world industrial food delivery recommendation system (Meituan) using user behavior logs. The dataset comprises three primary recommendation scenarios: Homepage (HP) Restaurant Recommendation, Shenqiangshou(SQS) Coupon-Package Recommendation , and Pinhaofan(PHF) Food Recommendation, with their respective data statistics summarized in Table 1. Specifically, HP serves as the highest-exposure scenario where restaurant cards are recommended to users for order placement, while the SQS and PHF scenarios are food-centric with distinct business objectives (detailed in 3.1.2). Over the past years, these scenarios have been continuously optimized both in business operations and recommendation architecture, making them the focal points for model innovation and industrial impact within Meituan’s ecosystem.
3.1.2 Metrics & Baselines
@@
To ensure rigorous comparison, we categorize baseline models according to their architecture types: 1) General recommenders including DCNv2 (Wang et al., 2021), MMoE (Ma et al., 2018a), and RankMixer (Zhu et al., 2025), which are widely adopted for feature interaction and multi-task learning. 2) Generative Ranking Models, such as MTGR (Han et al., 2025) and OneTrans (Zhang et al., 2025b), representing the latest advances in generative sequence modeling. 3) Multi-Scenario Recommendation Models, such as STAR (Sheng et al., 2021b) and PEPNet (Chang et al., 2023), which integrate multiple scenarios into a unified framework, allowing adaptive scenario modeling. Notably, models in (1) and (2) are trained on a single-scenario dataset, whereas (3) and (4) leverage a multi-scenario dataset for training.
3.1.3 Hyper-parameter Settings.
@@
| Generative Ranking | MTGR | 0.7679 | 0.6951 | 0.8776 | 0.6491 | 0.7883 | 0.7398 | 0.8879 | 0.7771 |
@@
| Our Method | MTFM | 0.7689 | 0.6954 | 0.8806 | 0.6507 | 0.7940 | 0.7474 | 0.8892 | 0.7824 |
| SQS | |||||||||
@@
| Generative Ranking | MTGR | 0.8454 | 0.7997 | 0.9097 | 0.8282 | 0.9095 | 0.8291 | 0.9059 | 0.8258 |
@@
| Our Method | MTFM | 0.8624 | 0.8027 | 0.9119 | 0.8301 | 0.9117 | 0.8319 | 0.9079 | 0.8288 |
3.2 Overall Performance Comparison (RQ1)
Table 2 presents the main experimental results for three representative scenarios: HP, PHF, and SQS. Several key insights can be drawn. First, RankMixer delivers competitive results among general recommenders by effectively supporting model scaling; it marginally exceeds other baselines on HP while achieving substantial gains on SQS, outperforming MMoE by 0.38pp in CTCVR GAUC and 0.49pp in WRITE GAUC. However, in the PHF setting, its performance remains superior to DCNv2 but trails behind MMoE. Second, multi-scenario methods such as STAR and PEPNet generally outperform general baselines but fall short of generative ranking models like MTGR and OneTrans. Notably, these methods exhibit a pronounced "see-saw effect"; for instance, PEPNet surpasses RankMixer in HP and PHF but underperforms in the SQS scenario. In contrast, generative models consistently demonstrate superior robustness, with MTGR slightly edging out OneTrans across most metrics. Ultimately, our proposed MTFM model achieves state-of-the-art results across nearly all scenarios and tasks. The consistent dominance of MTFM underscores its capability to not only mitigate the see-saw effect but, more importantly, to capitalize on the scaling law by efficiently leveraging extensive multi-scenario data to further enhance model's performance.
3.3 Efficiency Analysis of HTA (RQ2)
@@ 3.4 Scalability (RQ3)
3.4.1 Scalability with Model Size in MSR
@@
3.4.2 Scalability with Training Data Volumes in MRS
As depicted in Figure 4(b), we investigate the impact of the data scale on MTFM by selecting models of three different sizes and observing their performance fluctuations in SQS's CTCVR GAUC throughout the continual training phase. Empirical results reveal that across all model scales, performance improves consistently as the volume of training tokens increases. In experiments with larger data scales, we further observe that the performance gap between models of different sizes widens gradually as the number of training tokens increases. This robust scaling trend underscores MTFM’s data scalability and demonstrates its ability to consistently convert increased training volumes into performance improvements.
3.5 Interpretability Analysis (RQ4)
@@
3.6 Online Experiments (RQ5)
To validate MTFM's real-world impact, we conducted A/B tests across two major online scenarios. The experimental traffic involved tens of millions of daily exposures, ensuring robust results. The baseline model was the current SOTA model, which has been continuously optimized and deployed for several years.
@@ 3.6 Online Experiments (RQ5)
4 Related Work
4.1 Multi-Scenario Recommendation
Existing research primarily follows a "harmonize-then-decompose" paradigm, which aims to balance commonality and diversity by disentangling model parameters into domain-invariant and domain-specific components. For instance, STAR (Sheng et al., 2021a) pioneered this approach with a star topology that anchors domain-specific parameters around shared centered weights. To further capture complex feature interactions, M3OE (Zhang et al., 2024c) leverages three specialized MoE modules to learn hierarchical user preferences across common, domain, and task aspects. More recently, MLoRA (Yang et al., 2024) introduces a more parameter-efficient alternative by adapting LoRA modules for individual domains. Beyond this dominant paradigm, some studies (Zhang et al., 2023; 2024b) have explored graph-based modeling to represent heterogeneous information as distinct nodes. Nevertheless, these graph-based approaches often struggle to capture the information sufficiently between users and candidate items.
Moving beyond parameter-level disentanglement, recent studies have increasingly explored multi-scenario foundation models to achieve superior generalization. This line of work often adopts a "Foundation-Expert" architecture. For example, ExFM (Recommendation, 2025) distills knowledge from ultra-large-scale models into scenario-specific experts, while LFM4Ads (Zhang et al., 2025a) utilizes multi-granularity mechanisms to transfer comprehensive representations across features and tasks. Similarly, TAE(Li et al., 2025) incorporates HSTU as a robust encoder to bridge foundational target information with downstream expert models. However, these approaches predominantly rely on a two-stage pipeline rather than a fully end-to-end foundation model, which potentially limits the model's performance upper bound.
4.2 Ranking Model with Scalability
Current methods are primarily categorized into two paradigms: Transformer-based and non-Transformer-based (often referred to as Deep Learning Recommendation Models, DLRM). Transformer-based approaches partition input features into individual tokens and leverage Transformer-like architectures for unified modeling. Specifically, HSTU (Zhai et al., 2024) reconceptualizes recommendation as a sequential transduction task, employing a high-performance architecture to capture user behavior patterns. To further enrich feature interactions, MTGR (Han et al., 2025) extends this by incorporating cross-features and bidirectional attention mechanisms. Recognizing the computational demands of such models, OneTrans (Zhang et al., 2025b) addresses inference efficiency via a pyramid-style tail truncation strategy. More recently, scaling laws have been explored within generative frameworks (Deng et al., 2025; Zhou et al., 2025b), and these efforts also remain rooted in the Transformer backbone.
DLRM focuses mainly on feature interaction (Guo et al., 2017; Wang et al., 2021; Zhang et al., 2024a) and behavior sequence modeling (Zhou et al., 2018; 2019; Pi et al., 2020; Song et al., 2025) to enable ranking models to capture knowledge from various features (sequential/non-sequential). These approaches adopt a modular design, decomposing the system into distinct functional components and achieve scalability by selectively expanding specific stackable modules. For instance, RankMixer (Zhu et al., 2025) enhances the classic DLRM architecture by integrating Per-token Feed-Forward Networks to boost model capacity. Similarly, Wukong (Zhang et al., 2024a) attains scalability through a specialized design of stacked layers, combining Factorization Machine blocks with Linear Compress blocks to effectively increase model scalability.
5 Conclusion
In this paper, we propose MTFM, a recommendation foundation model that leverages heterogeneous tokenization to achieve a truly alignment-free architecture across diverse domains. Through the co-design of algorithms, training, and deployment systems, MTFM simultaneously embodies the three quintessential properties: Scalability, Extensibility, and Efficiency. Extensive offline evaluations demonstrate significant performance gains, establishing a new scalable paradigm that successfully transcends the limitations of single-scenario silos. Finally, online A/B tests confirm MTFM's real-world effectiveness, delivering substantial improvements in CTR and conversion volumes in Meituan.
6 Contributions
∗ Equal contribution; † Corresponding author. Contributors are listed alphabetically by their last names.
Core Contributors: Xin Song∗, Zhilin Guan∗, Ruidong Han∗†, Binghao Tang, Tianwen Chen, Bing Li, Zihao Li, Han Zhang, Fei Jiang†
Contributors: Chaolin Xie, Chi Ma, Chunyang Jiang, Chunzhen Jing, Dengxuan Li, Fengyi Li, Lei Yu, Mengyao Sun, Pu Wang, Qing Wang, Rui Fan, Shangyu Chen, Shifeng Du, Siyuan Bai, Wei Lin, Wentao Zhu, Zhou Han, Zhuo Chen, Zikang Xu

### v2

@@ Abstract.
1. Introduction
The landscape of Large Language Models (LLM) has been redefined by the era of Foundation Models (FMs), where unified architectures have evolved from unimodal to multimodal (Bai et al., 2025; Yang et al., 2025; Team et al., 2025; Fu et al., 2025; Liu et al., 2026). In parallel, the recommendation community has begun to validate the Scaling Laws (Kaplan et al., 2020; Hoffmann et al., 2022), primarily focusing on the expansion of model parameters and data volume within isolated scenarios (Zhai et al., 2024; Zhu et al., 2025; Zhang et al., 2024b; Wang et al., 2025; Zhou et al., 2025a; Han et al., 2024). However, the true potential of a Recommendation Foundation Model remains untapped as long as models are confined to single-scenario silos. Inspired by the success of multimodal FMs, we argue that the next frontier lies in bridging cross-scenario heterogeneity. By extending from single-scene optimization to a unified multi-scenario framework, we can leverage diverse behavioral signals to further unlock the benefits of Scaling Laws, thereby elevating the performance upper bound and establishing a more scalable paradigm for recommendation.
@@ 1. Introduction
Existing multi-scenario modeling efforts (Sheng et al., 2021a; Shu et al., 2024; Zhang et al., 2024a; Xiaoyu et al., 2025) primarily follow a ”harmonize-then-decompose” paradigm. These approaches typically aggregate data from multiple scenarios and enforce input homogeneity with a fixed template. Any heterogeneous components must be either discarded or handled through an additional padding operator. Structurally, they rely on disentangling model parameters into domain-invariant and domain-specific components (e.g., via MoE-based architectures (Ma et al., 2018a; Fedus et al., 2022)) to capture shared knowledge while preserving scenario-specific nuances.
@@
Empirical Validation: Extensive offline and online evaluations at Meituan demonstrate significant gains, proving MTFM’s real-world effectiveness.
2. Related Work
2.1. Multi-Scenario Recommendation
Existing research primarily follows a ”harmonize-then-decompose” paradigm, which aims to balance commonality and diversity by disentangling model parameters into domain-invariant and domain-specific components. For instance, STAR (Sheng et al., 2021a) pioneered this approach with a star topology that anchors domain-specific parameters around shared centered weights. To further capture complex feature interactions, M3OE (Zhang et al., 2024a) leverages three specialized MoE modules to learn hierarchical user preferences across common, domain, and task aspects. More recently, MLoRA (Yang et al., 2024) introduces a more parameter-efficient alternative by adapting LoRA modules for individual domains.
Moving beyond parameter-level disentanglement, recent studies have increasingly explored multi-scenario foundation models to achieve superior generalization. This line of work often adopts a ”Foundation-Expert” architecture. For example, ExFM (Recommendation, 2025) distills knowledge from ultra-large-scale models into scenario-specific experts, while LFM4Ads (Zhang et al., 2025b) utilizes multi-granularity mechanisms to transfer comprehensive representations across features and tasks. Similarly, (Li et al., 2025) incorporates HSTU as a robust encoder to bridge foundational target information with downstream expert models. However, these approaches predominantly rely on a two-stage pipeline rather than a fully end-to-end foundation model, which potentially limits the model’s performance upper bound.
2.2. Ranking Model with Scalability
Current methods are primarily categorized into two paradigms: Transformer-based and non-Transformer-based (often referred to as Deep Learning Recommendation Models, DLRM). Transformer-based approaches partition input features into individual tokens and leverage Transformer-like architectures for unified modeling. Specifically, HSTU (Zhai et al., 2024) reconceptualizes recommendation as a sequential transduction task, employing a high-performance architecture to capture user behavior patterns. To further enrich feature interactions, MTGR (Han et al., 2025) extends this by incorporating cross-features and bidirectional attention mechanisms. Recognizing the computational demands of such models, OneTrans (Zhang et al., 2025a) addresses inference efficiency via a pyramid-style tail truncation strategy. More recently, scaling laws have been explored within generative frameworks (Deng et al., 2025; Zhou et al., 2025b), and these efforts also remain rooted in the Transformer backbone. Conversely, traditional DLRM adopts a modular design, decomposing the system into distinct functional components. These methods achieve scalability by selectively expanding specific stackable modules. For instance, RankMixer (Zhu et al., 2025) enhances the classic DLRM architecture by integrating Per-token Feed-Forward Networks to boost model capacity. Similarly, WuKong (Zhang et al., 2024b) attains scalability through a specialized design of stacked layers, combining Factorization Machine blocks with Linear Compress blocks to effectively increase model scalability.
3. Methodology
@@
3.1. User Sample Aggregation with Multi-Scenario Data
@@
3.2. MTFM Model Architecture
@@ 3.2.1. Heterogeneous Tokenization
3.2.2. Hybrid Target Attention Architecture
@@
The embeddings of the T-tokens from the final layer are input into an MMoE(Ma et al., 2018a) module to compute prediction scores for various objectives across multiple scenarios.
@@
3.3. Training and Deployment Optimization
3.3.1. Data Pipeline of MTFM
@@ 4. Experiments
RQ1: How does the MTFM model perform across multiple heterogeneous multi-scenario tasks compared to other baseline models?
RQ2: How does the hybrid attention architecture in MTFM affect model training and inference performance?
RQ3: How does the MTFM model perform in terms of scalability, including model computational complexity scalability and data scalability?
@@
4.1. Experiments Setup
4.1.1. Datasets
@@
| HP | 240M | 4.23M | 18.53B | 1.08B | 176.77M |
| PHF | 151M | 8.07M | 15.29B | 359.14M | 104.73M |
| SQS | 44M | 0.98M | 2.24B | 85.34M | 9.92M |
To accommodate the heterogeneous multi-scenario business data proposed in this work, we construct our offline experimental dataset from a real-world industrial food delivery recommendation system (Meituan) using user behavior logs. The dataset comprises three primary recommendation scenarios: Homepage (HP) Restaurant Recommendation, Shenqiangshou(SQS) Coupon-Package Recommendation , and Pinhaofan(PHF) Food Recommendation, with their respective data statistics summarized in Table 1. Specifically, HP serves as the highest-exposure scenario where restaurant cards are recommended to users for order placement, while the SQS and PHF scenarios are food-centric with distinct business objectives (detailed in 4.1.2). Over the past years, these scenarios have been continuously optimized both in business operations and recommendation architecture, making them the focal points for model innovation and industrial impact within Meituan’s ecosystem.
4.1.2. Metrics & Baselines
@@
To ensure rigorous comparison, we categorize baseline models according to their architecture types: 1) General recommenders including DCNv2 (Wang et al., 2021), MMoE (Ma et al., 2018b), and RankMixer (Zhu et al., 2025), which are widely adopted for feature interaction and multi-task learning. 2) Generative Ranking Models, such as MTGR (Han et al., 2025) and OneTrans (Zhang et al., 2025a), representing the latest advances in generative sequence modeling. 3) Multi-Scenario Recommendation Models, such as STAR (Sheng et al., 2021b) and PEPNet (Chang et al., 2023), which integrate multiple scenarios into a unified framework, allowing adaptive scenario modeling. Notably, models in (1) and (2) are trained on a single-scenario dataset, whereas (3) and (4) leverage a multi-scenario dataset for training.
4.1.3. Hyper-parameter Settings.
@@
| Generative Ranking Method | MTGR | 0.7679 | 0.6951 | 0.8776 | 0.6491 | 0.7883 | 0.7398 | 0.8879 | 0.7771 |
@@
| Our Proposed Method | MTFM | 0.7689 | 0.6954 | 0.8806 | 0.6507 | 0.7940 | 0.7474 | 0.8892 | 0.7824 |
| Coupon-Package Recommendation (SQS) | |||||||||
@@
| Generative Ranking Method | MTGR | 0.8454 | 0.7997 | 0.9097 | 0.8282 | 0.9095 | 0.8291 | 0.9059 | 0.8258 |
@@
| Our Proposed Method | MTFM | 0.8624 | 0.8027 | 0.9119 | 0.8301 | 0.9117 | 0.8319 | 0.9079 | 0.8288 |
4.2. Overall Performance Comparison (RQ1)
Table 2 presents the main experimental results for three representative scenarios: HP, PHF, and SQS. Several key insights can be drawn. First, RankMixer delivers competitive results among general recommenders by effectively supporting model scaling; it marginally exceeds other baselines on HP while achieving substantial gains on SQS, outperforming MMoE by 0.38pp in CTCVR GAUC and 0.49pp in WRITE GAUC. However, in the PHF setting, its performance remains superior to DCNv2 but trails behind MMoE. Second, multi-scenario methods such as STAR and PEPNet generally outperform general baselines but fall short of generative ranking models like MTGR and OneTrans. Notably, these methods exhibit a pronounced ”see-saw effect”; for instance, PEPNet surpasses RankMixer in HP and PHF but underperforms in the SQS scenario. In contrast, generative models consistently demonstrate superior robustness, with MTGR slightly edging out OneTrans across most metrics. Ultimately, our proposed MTFM model achieves state-of-the-art results across nearly all scenarios and tasks. The consistent dominance of MTFM underscores its capability to not only mitigate the see-saw effect but, more importantly, to capitalize on the scaling law by efficiently leveraging extensive multi-scenario data to further enhance model’s performance.
4.3. Efficiency Analysis of HTA (RQ2)
@@ 4.4. Scalability (RQ3)
4.4.1. Scalability with Model Size in MSR
@@
4.4.2. Scalability with Training Data Volumes in MRS
As depicted in Figure 4(b), we investigate the impact of the data scale on MTFM by selecting models of three different sizes and observing their performance fluctuations in SQS’s CTCVR GAUC throughout the continual training phase. Empirical results reveal that across all model scales, performance improves consistently as the volume of training tokens increases. In experiments with larger data scales, we further observe that the performance gap between models of different sizes widens gradually as the number of training tokens increases. This robust scaling trend underscores MTFM’s data scalability and demonstrates its ability to consistently convert increased training volumes into performance improvements.
4.5. Interpretability Analysis (RQ4)
@@
4.6. Online Experiments (RQ5)
To validate MTFM’s real-world impact, we conducted A/B tests across two major online scenarios. The experimental traffic involved tens of millions of daily exposures, ensuring robust results. The baseline model was the current SOTA model, which has been continuously optimized and deployed for several years.
@@ 4.6. Online Experiments (RQ5)
5. Conclusion
In this paper, we propose MTFM, a recommendation foundation model that leverages heterogeneous tokenization to achieve a truly alignment-free architecture across diverse domains. Through the co-design of algorithms, training, and deployment systems, MTFM simultaneously embodies the three quintessential properties: Scalability, Extensibility, and Efficiency. Extensive offline evaluations demonstrate significant performance gains, establishing a new scalable paradigm that successfully transcends the limitations of single-scenario silos. Finally, online A/B tests confirm MTFM’s real-world effectiveness, delivering substantial improvements in CTR and conversion volumes in Meituan.